\begin{thebibliography}{10}
	%	\bibitem{ACI} Arkeryd, L., Cercignani, C., Illner, R.: Measure solutions of the steady Boltzmann equation in a slab, Comm. Math. Phys. {\bf 142}(2) 285-–296 (1991).
	\bibitem{AP} Anderson, J.L., Payne, A.C.: The relativistic Burnett equations and sound propagation. Physica A {\bf{85}} (1976) 261--286.
	        \bibitem{AW} Anderson, J.L., Witting, H.R.: A relativistic relaxational time model for the Boltzmann equation. Physica {\bf{74}} (1974) 466--488.
      \bibitem{BY} Bang, J.H., Yun, S.-B.: Stationary solution for the ellipsoidal BGK model in slab. J. Differential Equations. {\bf 261} (2016), 5803--5828.
        \bibitem{B} Bichteler, K.: On the Cauchy problem of the relativistic Boltzmann equation. Commun. Math. Phys. {\bf 4} (1967) 352--364.
		\bibitem{BCNS} Bellouquid, A., Calvo, J., Nieto, J., Soler, J.: On the relativistic BGK-Boltzmann model: asymptotics and hydrodynamics. J. Stat. Phys. {\bf 149} (2012) 284--316.
		\bibitem{BNU} Bellouquid, A., Nieto, J., Urrutia, L.: Global existence and asymptotic stability near equilibrium for the relativistic BGK model. Nonlinear Anal. {\bf 114} (2015) 87--104.
		%\bibitem{ABLP} Andries, P., Bourgat, J.-F.,  Le Tallec, P., Perthame, B.: Numerical comparison between the Boltzmann and
		%ES-BGK models for rarefied gases. Comput. Methods Appl. Mech. Engrg. {\bf 191} (2002), no. 31, 3369-3390.
		%\bibitem{ALPP} Andries,Pierre, Le Tallec, P., Perlat, J.-P., Perthame, B.: The Gaussian-BGK model of Boltzmann equation
		%with small Prandtl number. Eur. J. Mech. B Fluids {\bf 19} (2000), no. 6, 813-830.
		%\bibitem{ACI} Arkeryd, L., Cercignani, C., Illner, R. :  Measure solutions of the steady Boltzmann equation in a slab. Comm. Math. Phys. {\bf142} (1991), no. 2, 285–296.
		%\bibitem{AN1} Arkeryd, L.; Nouri, A. A compactness result related to the stationary Boltzmann equation in a slab, with applications to the existence theory. Indiana Univ. Math. J. {\bf44} (1995), no. 3, 815–839.
		%\bibitem{AN2} Arkeryd, L., Nouri, A. :  L1 solutions to the stationary Boltzmann equation in a slab. Ann. Fac. Sci. Toulouse Math.(6) {\bf9} (2000), no. 3, 375–413.
		%\bibitem{AN3} Arkeryd, L., Nouri, A. : The stationary Boltzmann equation in the slab with given weighted mass for hard and soft forces. Ann. Scuola Norm. Sup. Pisa Cl. Sci. (4) {\bf27} (1998), no. 3-4, 533–556 (1999).
		%\bibitem{B-G-L1} Bardos, C., Golse, F., Levermore, D.: Fluid dynamic limites of kinetic equations. I. Formal derivations,
		%J. Stat. Phys. {\bf 63} (1991), no. 1-2, 323-344.
		%\bibitem{B-G-L} Bardos, C., Golse, F., Levermore, D.: Fluid dynamic limites of kinetic equations. II. Convergence proofs for the Boltzmann equation,
		%Comm. Pure. Appl. Math  {\bf 46} (1993), 667-753.
		%\bibitem{A-G} Alonso, R.,  Gamba, I.M.: Propagation of $L^1$ and $L^{\infty}$ Maxweliian weighted bounds for derivatives of solutions
		%to solutions to the homogeneous elastic Boltzmann equation. J. Math. Pures. Appl. {\bf 89}, no.6, 575-595 (2007)
		%\bibitem{Ar1} Arkeryd, L.: $L^{\infty}$ estimates for the space-homogeneous Boltzmann equation, {\bf 31} no.2 347-361 (1983)
		%\bibitem{Ar2} Arkeryd, L.: On the Boltzmann equation, Arch. Rational. Mech. Anal. {\bf 45} 1-34 (1972)
		%\bibitem{A-E-P} Arkeryd, L. Esposito, R., Pulvirenti, M.: The Boltzmann equation for weakly inhomogeneous data. Comm. Math. Phys. {\bf 111} (3), 393-407 (1987)
		%\bibitem{Bello} Bellouquid, A.: Global existence and large-time behavior for BGK model for a gas with non-constant cross section, Transport Theory Statist. Phys. {\bf 32 } (2003) no. 2, 157-185.
	%	\bibitem{BGK} Bhatnagar, P. L., Gross, E. P. and Krook, M.: A model for collision processes in gases. Small amplitude process in charged
	%	and neutral one-component systems, Physical Revies, {\bf 94} (1954), 511-525.
		%\bibitem{B-P} Bouchut, F., Perthame, B.: A BGK model for small Prandtl number in the Navier-Stokes approximation.J. Stat. Phys. {\bf 71} (1993), no. 1-2, 191-207.
		%\bibitem{B1} Brull, S.: The stationary Boltzmann equation for a two-component gas for soft forces in the slab. Math. Methods Appl. Sci. {\bf31} (14) (2008) 1653--1666, 35F30 (76P05 82C40)
		%\bibitem{B2} Brull, S.: The stationary Boltzmann equation for a two-component gas in the slab with different molecular masses. Adv. Differential Equations {\bf15} (11--12) (2010) 1103--1124
		%\bibitem{B3} Brull, S., Scneider, J.: A new approach for the ellipsoidal statistical model. Contin. Mech. Thermodyn. {\bf20} (2008), no.2, 63-74.
		%\bibitem{B2} Brull, S., Schneider, J.: Derivation of a BGK model for reacting gas mixtures. Commun. Math. Sci. {\bf 12} (2014), 117-125.
		%\bibitem{Car}Carleman, T.: Problems mathematiques dans la theorie cineqitue des gas Uppsala: Almquist \& Wiksells 1957
		%\bibitem{Caf1}  Caflisch, Russel E. The Boltzmann equation with a soft potential. I. Linear, spatially-homogeneous. Comm. Math. Phys. 74 (1980), no. 1, 71–95.
		%\bibitem{Caf2} Caflisch, Russel E. The Boltzmann equation with a soft potential. II. Nonlinear, spatially-periodic. Comm. Math. Phys. 74 (1980), no. 2, 97–109.
		\bibitem{Cal} Calogero, S.: The Newtonian limit of the relativistic Boltzmann equation. J. Math. Phys. {\bf45} (2004) 4042--4052.
		%\bibitem{C} Cercignani, C.: The Boltzmann Equation and Its Application. Springer-Verlag, 1988.
		%\bibitem{CIS} Cercignani, C., Illner, R., Shinbrot, M. : A boundary value problem for discrete-velocity models. Duke Math. J. 55 (1987), no. 4, 889–900.
		%\bibitem{CC} Chapman, C. and Cowling, T. G.: The mathematical theory of non-uniform gases, Cambridge University Press, 1970.
		%\bibitem{CIP} Cercignani, C., Illner, R., Pulvirenti, M.: The Mathematical Theory of Dilute Gases. Springer-Verlag, 1994.
		\bibitem{CK} Cercignani, C., Kremer, G.M.: The Relativistic Boltzmann Equation: Theory and Applications. Birkh\"{a}user, Berlin (2003).
		\bibitem{CK2} Cercignani, C., Kremer, G.M.: Moment closure of the relativistic Anderson
		and Witting model equation. Physica A {\bf{290}} (2001) 192--202.
		\bibitem{CKT} Chen, Y., Kuang, Y., Tang, H.: Second-order accurate genuine BGK schemes for the ultra-relativistic flow simulations. J. Comput. Phys. {\bf{349}} (2017), 300--327.
		%\bibitem{Chan} Chan, W. M.: An energy method for the BGK model. M. Phil thesis, City University of Hong Kong, 2007.
		%\bibitem{De Groot} De Groot, S.R., van Leuwen, W.A., van Weert, Ch.G.: Relativistic Kinetic Theory.
		%North Holland, Amsterdam (1980)
		%\bibitem{Duan} Duan, R.: hypocoercivity of the linearized dissipative kinetic equations, Nonlinearity, {\bf 24} (2011), no. 8, 2165-2189
        \bibitem{D} Dudy\'{n}ski, M.: On the linearized relativistic Boltzmann equation. II. Existence of hydro-dynamics.
        J. Stat. Phys. {\bf 57} (1-–2) (1989) 199--245.	
        \bibitem{DE} Dudy\'{n}ski, M., Ekiel-Je\.{z}ewska, M.L.: On the linearized relativistic Boltzmann equation. I. Existence of
        solutions. Commun. Math. Phys. {\bf 115} (4) (1985) 607--629.
		%\bibitem{Duan-StrainV} Duan, R, Strain, R.: Optimal large-time behavior of the Vlasov-Maxwell-Boltzmann system in the whole space. Comm. Pure. Appl. Math. {\bf 64} (2011),
		%no2, 375-413.
		%\bibitem{Duan-STrainM} Duan, R, Strain, R.: Optimal time decay of the Vlasov-Poisson-Boltzmann system in $\mathbb{R}^3$. Arch. Rational. Mech. Anal. {\bf 199}(2010), no.1, 291-328.
		%\bibitem{D-U-Y} Duan, R., Ukai, S., Yang, Tong; Zhao, H.: Optimal decay estimates on the linearized Boltzmann equation with time dependent force and their applications.Comm. Math. Phys. {\bf277} (2008), no. 1, 189–236.
		%\bibitem{D-L1} Diperna, R., Lions, P.-L.: On the Cauchy problem for the Boltzmann equation: Global exitence and weak stability. Ann. Math.
		%{\bf130} (2), 321-366 (1989).
		%\bibitem{G-P-V} Gamba, I.M., Panferov, V., Villani, C : Upper Maxwellian Bounds for the spatially homogeneous Boltzmann equation.
		%Arch. Rational Mech. Anal. {\bf 194} (2009) 253-282.
		%\bibitem{Dud1} Dudyński, M. : On the linearized relativistic Boltzmann equation. II. Existence of hydrodynamics, J. Stat. Phys. 57 (1989) 199–245.
		%\bibitem{Dud2} Dudyński, M., Ekiel-Jeżewska, M. L. : On the linearized relativistic Boltzmann equation. I. Existence of solutions, Comm. Math. Phys. 115 (1988) 607–62
		\bibitem{Dud3} ] Dudyński, M., Ekiel-Jeżewska, M.L.: Global existence proof for relativistic Boltzmann equation, J. Stat. Phys. {\bf 66} (3) (1992) 991--1001.
	  %  \bibitem{Einstein1} Einstein, A.: \"{U}ber einen die Erzeugung und Verwandlung des Lichtes betreffenden heuristischen Gesichtspunkt. Annalen der Physik {\bf{17}} (6), 132-148 (1905).
	%	\bibitem{Einstein2} Einstein, A.: \"{U}ber die von der molekularkinetischen Theorie der W\"{a}rme geforderte Bewegung von in ruhenden Fl\"{u}ssigkeiten suspendierten Teilchen. Annalen der Physik. {\bf{17}} (8), 549–560 (1905).
	%	\bibitem{Einstein3} Einstein, A,: Zur Elektrodynamik bewegter K\"{o}rper. Annalen der Physik. {\bf{17}} (10), 891–921 (1905).
	%	\bibitem{Einstein4} Einstein, A.: Ist die Tr\"{a}gheit eines K\"{o}rpers von seinem Energieinhalt abh\"{a}ngig?. Annalen der Physik. {\bf{18}} (13): 639–641 (1905).
		%\bibitem{F-J} Filbet, F., Jin, S.: An asymptotic preserving scheme for the ES-BGK model of the Boltzmann equation. J. Sci. Comput. {\bf 46} (2011), no.2, 204-224.
		%\bibitem{F-R} Filbet, F., Russo, G.: Semilagrangian schemes applied to moving boundary problems for the BGK model of rarefied gas dynamics, Kinet. Relat.
		%Models {\bf 2} (2009), no.1, 231-250.
		\bibitem{FRS} Florkowski, W., Ryblewski, R., Strickland, M.: Anisotropic hydrodynamics for rapidly expanding systems. Nuclear Physics A {\bf{916}} (2013) 249--259.
		%\bibitem{G-T} Galli, M.A., Torczynski, R.: Investigation of the ellipsoidal-statistical Bhatnagar-Gross-Krook kinetic model applied to gas-phase transport of heat and tangential momentum between parallel walls, Phys. Fluids, {\bf 23} (2011) 030601
		%\bibitem{GL} Glassey, R.: The Cauchy Problmes in Kinetic Theory. SIAM 1996.
		%Glassey, Robert T.(1-IN)
		%\bibitem{Glassey V} Glassey, R. : Global solutions to the Cauchy problem for the relativistic Boltzmann equation with near-vacuum data. Comm. Math. Phys. 264 (2006), no. 3, 705–724.
		\bibitem{GS1} Glassey, R.T., Strauss, W.A.: Asymptotic stability of the relativistic Maxwellian. Publ. Res. Inst. Math. Sci. {\bf 29} (2) (1993) 301--347.
		\bibitem{GS2} Glassey, R.T., Strauss, W.A.: Asymptotic stability of the relativistic Maxwellian via fourteen moments. Trans. Th. Stat. Phys. {\bf 24} (4--5) (1995) 657--678.
		%\bibitem{GS3}  Glassey, R. T., Strauss, W. A. :  On the derivatives of the collision map of relativistic particles. Transport Theory Statist. Phys. 20 (1991), no. 1, 55–68.
		%\bibitem{GL3} Glassey, R. T., Strauss, W. A. : The relativistic Boltzmann equation near equilibrium. Spectral and scattering theory and applications, 105–111, Adv. Stud. Pure Math., 23, Math. Soc. Japan, Tokyo, 1994.
		
		%\bibitem{Gome}  Ghomeshi, S. :  Existence and uniqueness of solutions for the Couette problem. J. Stat. Phys. {\bf118} (2005), no. 1-2, 265–300.
		%\bibitem{Guo whole} Guo, Y.: The Boltzmann equation in the whole space. Indiana Univ. Math. J. {\bf 53} (2004). no.4, 1081-1094
		%\bibitem{Guo VMB} Guo, Y.: The Vlasov-Maxwell-Boltzmann system near Maxwellians. Invent. Math. {\bf 153} (2003) no.3, 593-630
		%\bibitem{Guo VPB} Guo, Y.: The Vlasov-Poisson-Boltzmann system near Maxwellians. Comm. Pure. Appl. Math., {\bf 55} (2002) no.9, 1104-1135.
		%\bibitem{Guo CD} Guo, Y.: Continuity and decay of the Boltzmann equation in bounded domains. Arch. Ration. Mech. Anal. {\bf 197} (2010) no.3, 713-809 (2010)
		%\bibitem{Guo Strain} Guo, Y. Strain, R.: Almost exponential decay near Maxwellian, Comm. P.D.E. {\bf 31} (2006)no.2,  543-567.
		\bibitem{Guo Strain Momentum} Guo, Y., Strain, R.M.: Momentum regularity and stability of the relativistic Vlasov-Maxwell-Boltzmann system. Comm. Math. Phys. {\bf 310} (3) (2012) 649--673.
		
		%\bibitem{Guo-Wang} Guo, Y., Wang, Y.: Decay of dissipative equations and negative Sobolev spaces. Comm. Partial Differential Equations {\bf 37} (2012), no. 12, 2165–2208.
		
		\bibitem{HM} Hakim, R., Mornas, L.: Collective effects on transport coefficients of relativistic nuclear matter. Phys. Rev. C {\bf{47}} (1993) 2846.
		
		\bibitem{HMPS} Hakim, R., Mornas, L., Peter, P., Sivak, H.D.: Relaxation time approximation for relativistic dense matter. Phys. Rev. D {\bf{46}} (1992) 4603.
		%\bibitem{Holway} Holway, L.H.: Kinetic theory of schock structure using and ellipsoidal distribution function. Rarefied Gas Dynamics, Vol. I (Proc. Fourth Internat. Sympos., Univ. Toronto, 1964), Academic Press, New York, (1966), pp. 193-215.
		%\bibitem{Issau} Issautier, D.: Convergence of a weighted particle method for solving the Boltzmann (B.G,K.) equaiton, Siam Journal on Numerical Analysis,{\bf 33}, no 6 (1996), 2099-2199.
		%\bibitem{Ha Strain} Ha, S.-Y., Jeong, E., Strain, R.: Uniform $L^1$-stability of the relativistic Boltzmann equation near vacuum.  Commun. Pure Appl. Anal. 12 (2013), no. 2, 1141–1161.
        \bibitem{Jut1} J\"{u}ttner, F.: Das Maxwellsche Gesetz der Geschwindigkeitsverteilung in der Relativtheorie.
		Ann. Phys. {\bf 339} (5) (1911) 856--882.
		%\bibitem{Jut2} 38. J\"{u}ttner, F.: Die relativische Quantentheorie des idealen Gases. Z. Phys. 47, 542–566 (1928)
		\bibitem{Jiang1} Jiang, Z.: On the relativistic Boltzmann equation, Acta Math. Sci. (Engl. Ed.) {\bf 18} (3) (1998) 348--360.
		\bibitem{Jiang2} Jiang, Z: On the Cauchy problem for the relativistic Boltzmann equation in a periodic box: global existence. Transport Theory Statist. Phys. {\bf 28} (6) (1999) 617--628.
		\bibitem{LM} Lichnerowicz, A., Marrot, R.: Propriétés statistiques des ensembles de particules en relativité restreinte. C. R. Acad. Sci. Paris {\bf 210} (1940) 759--761.
		\bibitem{LR} Lee, H., Rendall, A.: The spatially homogeneous relativistic Boltzmann equation with a hard potential, Comm. Partial Differential Equations {\bf 38} (12) (2013) 2238--2262.
		%\bibitem{Kawashima} Kawashima, S. : The Boltzmann equation and thirteen moments, Japan J. Appl. Math. {bf 7} (1990), 301-320.
		\bibitem{Kremer} Kremer, G. M.: Diffusion of relativistic gas mixtures in gravitational fields. Physica A {\bf{393}} (2014) 76--85.
        \bibitem{Kremer2} Kremer, G. M.: Relativistic gas in a Schwarzschild metric. J. Stat. Mech. (2013) P04016.
		%\bibitem{Maslova} Maslova, N.B., Tchubenko, R.P.: Dokl. Akad. Nauk. SSSR {\bf 202} (1973), 800-803. Vesin. Leningr. Univ. {\bf 1} (1973), 100-105.
		%{\bf 7}, 109-113 (1976). Vestn. Leningr. Univ. {\bf 13}, 90-97 (1976)
		%\bibitem{Mouhot} Mouhot, C.: Rate of convergence to equilibrium for the spatially homogeneous Boltzmann equation with hard potentials.
		%Comm. Math. Phys. {\bf 261} no.3 629-672 (2006)
		%\bibitem{L-Y-Y} Liu, T.-P, Yang, T. Yu, S.-H.: Energy method for Boltzmann equation. Phys. D {\bf 188} (2004), no. 3-4, 178-192.
		%\bibitem{L-Y1} Liu, T.-P.; Yu, S.-H. : The Green's function and large-time behavior of solutions for the one-dimensional Boltzmann equation.Comm. Pure Appl. Math. {\bf57} (2004), no. 12, 1543–1608.
		%\bibitem{L-Y2} Liu, T.-P.; Yu, S.-H. : Solving Boltzmann equation, Part I: Green's function. Bull. Inst. Math. Acad. Sin. (N.S.) {\bf6} (2011), no. 2, 115–243.
		\bibitem{Majorana} Majorana, A.: Relativistic relaxation models for a simple gas. J. Math. Phys. {\bf{31}} (8) (1990) 2042--2046.
		\bibitem{Mar1} Marle, C.: Sur l\textquoteright\'{e}tablissement des equations de l\textquoteright hydrodynamique des fluides relativistes dissipatifs, I. L\textquoteright equation de Boltzmann relativiste. Ann. Inst. Henri Poincar\'{e} {\bf{10}} (1969) 67--127.
		\bibitem{Mar2} Marle, C.: Sur l\textquoteright\'{e}tablissement des equations de l\textquoteright hydrodynamique des fluides relativistes dissipatifs. II. M\'{e}thodes de r\'{e}solution approch\'{e}e de l\textquoteright equation de Boltzmann
		relativiste. Ann. Inst. Henri Poincar\'{e} {\bf 10} (1969) 127--194.
		\bibitem{Mar3} Marle, C.: Modele cin\'{e}tique pour l\textquoteright \'{e}tablissement des lois de la conduction de la
		chaleur et de la viscosit\'{e} en th\'{e}orie de la relativit\'{e}. C. R. Acad. Sci. Paris {\bf{260}} (1965) 6539--6541.
		\bibitem{MBHS} Mendoza, M., Boghosian, B.M., Herrmann, H.J., Succi, S.: Derivation of the lattice Boltzmann model for relativistic hydrodynamics. Phys. Rev. D {\bf{82}} (2010) 105008.
		
		
		%\bibitem{M-M-M} Mellet, A., Mischler, S., Mouhot, C.: Fractional diffusion limit for collisional kinetic equations. Arch. Ration. Mech. Anal. 199 (2011), no. 2, 493–525.
		%\bibitem{Mellet} Mellet, A.: Fractional diffusion limit for collisional kinetic equations: a moments method. Indiana Univ. Math. J. 59 (2010), no. 4, 1333–1360.
		%\bibitem{M-S} Mieussens, L., Struchtrup, H. : Numerical comparison of Bhatnagar-Gross-Krook models with proper Prandtl number, Phys. Fluids {\bf 16}(2004), no.8. 2797-2813{\bf 10} (2000), no. 8, 1121-1149.
		%\bibitem{M-W-R}Meng, J., Wu, L., Reese, J. M., Zhang, Y.: Assessment of the ellipsoidal-statistical Bhatnagar-Gross-Krook model for force-driven Poiseuille flows. J. Comput. Phys. 251 (2013), 383–395.
		%\bibitem{Mischler} Mischler, S.: Uniqueness for the BGK-equation in $R^n$ and the rate of convergence for a semi-discrete scheme.
		%Differential integral Equations {\bf 9} (1996), no.5, 1119-1138.
		%\bibitem{Perthame} Perthame, B. : Global existence to the BGK model of Boltzmann equation. J. Differential Equations. {\bf 82} (1989), no.1, 191-205.
		%\bibitem{P-Puppo} Pieraccini, S., Puppo, G.: Implicit-explicit schemes for BGK kinetic equations. J. Sci. Comput. {\bf 32} (2007), no.1, 1-28.
		%\bibitem{P-P} Perthame, B., Pulvirenti, M. : Weighted $L^{\infty}$ bounds and uniqueness for the Boltzmann BGK model. Arch. Rational Mech. Anal. {\bf 125}
		%(1993), no. 3, 289-295.
		%\bibitem{R-S-Y} Russo, G., Santagati, P. and Yun, S.-B. : Convergence of a semi-Lagrangian scheme for the BGK model of the Boltzmann equation. arXiv: 1007.2843v1 [math.AP].
		%\bibitem{SR1} Saint-Raymond, L.: From the BGK model to the Navier-Stokes equations. Ann. Sci. Ecole Norm. Sup {\bf 36} (2003), no.2, 271-317.
		%\bibitem{SR1} Saint-Raymond, L.: Discrete time Navier-Stokes limit for the BGK Boltzmann equation. Comm. Partial Differential Equations {\bf 27} (2002), no. 1-2,149-184.
		%\bibitem{SR2} Saint-Raymond, L.: From the BGK model to the Navier-Stokes equations. Ann. Sci. Ecole Norm. Sup {\bf 36} (2003), no.2, 271-317.
		%\bibitem{Sone} Sone, Y.: Kinetic Theory and Fluid Mechanics. Boston: Birkh\"{a}user, 2002.
		%\bibitem{Sone2} Sone, Y.: Molecular Gas Dynamics: Theory, Techniques, and Applications. Boston: Brikh\"{a}user, 2006.
		\bibitem{SS} Speck, J., Strain, R.M.: Hilbert expansion from the Boltzmann equation to relativistic fluids. Commun. Math. Phys. {\bf 304} (2011) 229--280.
		\bibitem{Strain1} Strain, R.M.: Asymptotic stability of the relativistic Boltzmann equation for the soft potentials. Commun. Math. Phys. {\bf 300} (2010) 529--597.
		%\bibitem{Strain coordinate}  Strain, M. :  Coordinates in the relativistic Boltzmann theory. Kinet. Relat. Models 4 (2011), no. 1, 345–359.
		\bibitem{Strain2} Strain, R.M.: Global Newtonian limit for the relativistic Boltzmann equation near vacuum. SIAM J. Math. Anal. {\bf 42} (2010) 1568--1601.
		%\bibitem{Strain Guo} Strain, Robert M.; Guo, Yan Exponential decay for soft potentials near Maxwellian. Arch. Ration. Mech. Anal. 187 (2008), no. 2, 287–339.
		%\bibitem{Strain Guo2}    Strain, R., Guo, Y. : Stability of the relativistic Maxwellian in a collisional plasma.
		%Comm. Math. Phys. 251 (2004), no. 2, 263–320.
        \bibitem{Strain Yun} Strain, R.M., Yun, S.-B.: Spatially homogenous Boltzmann equation for relativistic particles, SIAM J. Math. Anal. {\bf 46} (1) (2014) 917--938.
        \bibitem{Strain Zhu} Strain, R.M., Zhu, K.: Large-time decay of the soft potential relativistic Boltzmann equation in $\mathbb{R}^3_x$. Kinet. Relat. Models {\bf 5} (2) (2012) 383--415.
		
		%60. Ukai, S.: On the existence of global solutions of mixed problem for
		%\bibitem{Strain}  Strain, R. M.: The Vlasov-Maxwell-Boltzmann system in the whole space. Comm. Math. Phys. {\bf268} (2006), no. 2,543–567.
		%\bibitem{Stru} Struchtrup, H.: The BGK-model with velocity-dependent collision frequency. Contin. Mech. Thermodyn. {\bf 9} (1997), no.1 , 23-31.
        \bibitem{Stru2} Struchtrup, H.: Projected moments in relativistic kinetic theory. Physica A {\bf{253}} (1998) 555--593.
		
		%\bibitem{Stru-book} Struchtrup, H.: Mesoscopic transport equaitons for rarefied gas flows: Approximation methods in kinetic theory. Springer. 2005.
        \bibitem{TH} Tenreiro, R.D., Hakim, R.: Transport properties of the relativistic degenerate electron gas in a strong magnetic field: Covariant relaxation-time model. Phys. Rev. D {\bf{15}} (1977) 1435.
        \bibitem{TI} Takamoto, M., Inutsuka, S.-I.: The relativistic kinetic dispersion relation: Comparison of the relativistic Bhatnagar-Gross-Krook model and Grad's 14-moment expansion. Physica A {\bf{389}} (2010) 4580--4603.
	%	\bibitem{Ukai} Ukai, S.: On the existence of global solutions of a mixed problem for the nonlinear Boltzmann equation. Proc. Japan Acad., Ser. A
		%{\bf 53}, 179-184 (1974)
		%\bibitem{Ukai} Ukai, S.: On the existence of global solutions of mixed problem for non-linear Boltzmann equation. Proc. Japan Acad. {\bf50} (1974), 179–184.
		%\bibitem{Ukai-BGK} Ukai, S.: Stationary solutions of the BGK model equation on a finite interval with large boundary data. Transport theory Statist. Phys. {\bf 21} (4--6) (1992)
		%\bibitem{U-T} Ukai, S. Yang, T.: Mathematical Theory of Boltzmann equation, Lecture Notes Series. no. 8, Liu Bie Ju Center for Math. Sci,City University of Hong Kong, 2006.
		%\bibitem{U-T1} Ukai, Seiji; Yang, Tong The Boltzmann equation in the space L2∩L∞β: global and time-periodic solutions. Anal. Appl. %(Singap.) 4 (2006), no. 3, 263–310.
		%\bibitem{YY} Yang, T., Yu, H. : Hypocoercivity of the relativistic Boltzmann and Landau equations in the whole space
		%J. Differential Equations 248 (2010), no. 6, 1518–1560.
		%\bibitem{V} Villani, C.: A Review of mathematical topics in collisional kinetic theory. Handbook of mathematical fluid dynamics. Vol. I. North-Holland. Amsterdam, 2002, 71-305
		%\bibitem{Wal} Walender, P.: On the temperature jump in a rarefied gas, Ark, Fys. {\bf 7}  (1954), 507-553.
		%\bibitem{Yang Yu1} Yang, T., Yu, H.-J.: Hypocoercivity of the relativistic Boltzmann and Landau equations in the whole space. J. Diff. Eqs. {\bf248} (6), 1518–1560 (2010)
		%\bibitem{Yang Yu2} Yang, T., Yu, H.-J.: Optimal Convergence Rates of Classical Solutions for Vlasov-Poisson-Boltzmann System,
		%Comm. Math. Phys. {\bf 301} 319-355 (2011)
		%\bibitem{Yun1} Yun, S.-B.: Cauchy problem for the Boltzmann-BGK model near a global Maxwellian. J. Math. Phy. {\bf 51} (2010), no. 12, 123514, 24pp.
		%\bibitem{Yun2} Yun, S.-B.: Classical solutions for the ellipsoidal BGK model with fixed collision frequency.
		% J. Differential Equations  259  (2015),  no. 11, 6009–6037
		%\bibitem{Yun3} Yun, S.-B.: Ellipsoidal BGK Model Near a Global Maxwellian. SIAM J. Math. Anal. {\bf47} (2015), no. 3, 2324–2354.
		%\bibitem{Zhang} Zhang, X.: On the Cauchy problem of the Vlasov-Posson-BGK system: global existence of weak solutions. J. Stat. Phys. {\bf141} (2010),no.3, %566-588.
		%\bibitem{Z-H} Zhang, X., Hu, S.: $L^p$ solutions to the Cauchy problem of the BGK equation. J. Math. Phys. {\bf 48} (2007) no.11, 113304, 17pp.
		%\bibitem{Z-VPBGK} Zhang, X.: On the Cauchy problem of the Vlasov-Poisson-BGK system: global existence of weak solutions. J. Stat. Phys. 141 (2010), no. 3, 566–588
		%\bibitem{Z-Stru} Zheng, Y., Struchtrup, H. : Ellipsoidal statistical Bhatnagar-Gross-Krook model with velocity dependent collision frequency. Phys. Fluids {\bf 17} (2005), 127103, 17pp.
	\end{thebibliography}

\documentclass[10pt]{amsart}
\usepackage[margin=1.5in]{geometry}
\usepackage{color}
\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{proposition}{Proposition}[section]

\theoremstyle{definition}
\newtheorem{definition}[theorem]{Definition}

\theoremstyle{remark}
\newtheorem{remark}[theorem]{Remark}

\numberwithin{equation}{section}

\usepackage{graphicx}
\DeclareGraphicsExtensions{.pdf,.png,.jpg}
\begin{document}
	\title[Staionary solution for relativistic BGK model]{Stationary solutions to the boundary value problem for relativistic BGK model in a slab}
	\author[B.-H. Hwang]{Byung-Hoon Hwang}
	\address{Department of Mathematics, Sungkyunkwan University, Suwon 440-746, Republic of Korea}
	\email{bhh0116@skku.edu}

	\author[S.-B. Yun]{Seok-Bae Yun}
	\address{Department of Mathematics, Sungkyunkwan University, Suwon 440-746, Republic of Korea}

	\email{sbyun01@skku.edu}

	\keywords{relativistic BGK model, relativistic Boltzmann equation, boundary value problem, stationary solution, kinetic theory of gases}
	
\begin{abstract}
 In this paper, we are concerned with the boundary value problem in a slab for the stationary relativistic BGK model of Marle type, which  is a relaxation model of the relativistic Boltzmann equation.
In the case of fixed inflow boundary conditions, we  establish the existence of unique stationary solutions.
 \end{abstract}
	\maketitle

\section{Introduction}
In this paper, we address the existence of stationary solutions for a relativistic BGK model defined on a unit interval:
\begin{equation}\label{SRBGK}
	q_1\frac{\partial f}{\partial x}=w\big(J_f-f\big),\quad (x,q)\in [0,1]\times \mathbb{R}^3,
\end{equation}
endowed with a fixed inflow data at the boundary:
$$
f(0,q)=f_L(q)~~ \text{for}~~ q_1>0,\qquad f(1,q)=f_R(q)~~ \text{for}~~ q_1<0,
$$
for some given functions $f_L$ and $f_R$.
The momentum distribution function $f(x,q)$ represents the number density of relativistic particles at position $x\in [0,1]$ with momentum $q\in \mathbb{R}^3$. On the r.h.s of (\ref{SRBGK}), $w$ is a collision frequency, and $J_f$ denotes the local relativistic Maxwellian defined by
$$
J_f=\frac{n}{M(\beta)}e^{-\beta\left( \sqrt{1+|u|^2}\sqrt{1+|q|^2}-u\cdot q\right)},
$$
where $M(\beta)$ is
$$
M(\beta)=\int_{\mathbb{R}^3}e^{-\beta\sqrt{1+|p|^2}}dp,
$$
and the proper particle density $n$, velocity four-vector $(\sqrt{1+|u|^2},u)$ and the equilibrium temperature $1/\beta$ are defined by the following relations:
(in the following, $q_0$ denotes $\sqrt{1+|q|^2}$ for  $q\in\mathbb{R}^3$).
\begin{align}\label{macro fields}
\begin{split}
 	&\hspace{0.2cm}n^2=\biggl(\int_{\mathbb{R}^3}f dq\biggl)^2-\sum_{i=1}^3\biggl(\int_{\mathbb{R}^3}f\frac{q_i}{q_0} dq\biggl)^2,\cr
 	& n\sqrt{1+|u|^2}=\int_{\mathbb{R}^3}fdq,\quad n u=\int_{\mathbb{R}^3}f\frac{q}{q_0}dq,\cr
 	&\hspace{1.3cm}\frac{K_1}{K_2}(\beta)=\frac{1}{n}\int_{\mathbb{R}^3}f \frac{1}{q_0}dq,
\end{split}
\end{align}
where $K_i$ denotes a modified Bessel function of the second kind:
\[
K_i(\beta)=\int_0^\infty \cosh(ir)e^{-\beta\cosh(r)}dr \quad(i=1,2).
\]
It is shown in  \cite{BCNS} that $K_1/K_2$ is strictly increasing and, therefore,
the last identity in (\ref{macro fields}) uniquely determines $\beta$.

The kinetic theory of relativistic particles began with J\"{u}ttner \cite{Jut1} in 1911 when he derived a relativistic version of the Maxwellian distribution, which is often called the J\"{u}ttner equilibrium. The relativistic generalization of the celebrated Boltzmann equation was made by Lichnerowicz and Marrot in 1941 \cite{LM}.

The complicated structure of the relativistic collision operator, however, has long been a major obstacle in the application of the relativistic Boltzmann equation to various flow problems. To circumvent this difficulty, two types of relaxation in time approximation were suggested \cite{AW,Mar2, Mar3} to develop a numerically amenable model equation which still shares essential features of the collision operator such as the conservation laws and H-theorem. The first one was proposed by Marle \cite{Mar2, Mar3} where the macroscopic fields are represented using the Eckart decomposition, and the other one by Anderson and Witting \cite{AW} where the Landau-Lifshitz decomposition was employed for the representation of the macrosocipic fields.


The relativistic BGK models then have been widely used for various purposes \cite{AP,AW,CK,CK2,CKT,FRS,HM,HMPS,Kremer,Kremer2,Majorana,Mar1,Mar3,MBHS,Stru2,TH,TI}, but rigorous mathematical studies have just started and lots of issues still remain to be addressed. In 2012, Bellouquid et al considered the determination of equilibrium parameters, various formal scaling limits  and the analysis of the linearized problem of the Marle model in \cite{BCNS}. Then, the existence of mild solutions and asymptotic stability of the Marle model near global relativistic equilibrium were proved  in \cite{BNU}. To the best knowledge of authors, these two works are the only mathematical literatures treating the relativistic BGK model analytically.

Much more have been done for the relativistic Boltzmann equation. We refer to \cite{B} for local existence, and \cite{D,DE,GS1,GS2,Strain1,Strain Zhu} for global existence near equilibrium. Momentum regularity was established in  \cite{Guo Strain Momentum} leading to the global existence for the relativistic Vlasov-Maxwell-Boltzmann equation near equilibriums. For the existence of renormalized solutions with general data, see \cite{Dud3,Jiang1,Jiang2}. Existence and some moment estimates for the spatially homogeneous relativistic Boltzmann equation can be found in \cite{LR,Strain Yun}. We also refer to \cite{Cal,Strain2} for Newtonian limits and \cite{SS} for hydrodynamic limits.



\subsection{Main result}
In this paper, we consider the stationary relativistic BGK model of Marle type posed on a bounded interval with fixed inflow boundary data at both ends, and establish the existence of unique stationary solutions.
\begin{definition}\label{mild}
A non-negative function $f\in L^1\left([0,1]\times\mathbb{R}^3\right)$ is called a mild solution of \eqref{SRBGK} if
\begin{align*}
f(x,q)&=\left(e^{-\frac{w}{|q_{1}|}x}f_L+\frac{w}{|q_1|}\int_0^xe^{-\frac{w}{|q_1|}(x-y)}J_fdy\right)1_{q_1>0}\cr
&+\left(e^{-\frac{w}{|q_{1}|}(1-x)}f_R+\frac{w}{|q_1|}\int_x^1e^{-\frac{w}{|q_1|}(y-x)}J_fdy\right)1_{q_1<0}.
\end{align*}
\end{definition}
\noindent For brevity, we denote
\begin{align*}
f_{LR}&=f_L1_{q_1>0}+f_R1_{q_1<0},\cr
f^e_{LR}&=e^{-\frac{w}{|q_{1}|}x}f_{L}1_{q_1>0}+e^{-\frac{w}{|q_{1}|}(1-x)}f_{R}1_{q_1<0}.
\end{align*}
We then define quantities $a_l, a_u$ and $\lambda$ by
\begin{align*}
a_l=\int_{\mathbb{R}^3} e^{-\frac{w}{|q_1|}}f_{LR}\frac{1}{q_0^2} dq,\qquad a_u=2\int_{\mathbb{R}^3} f_{LR}dq,
\end{align*}
and
\begin{align}\label{lambda}
	\lambda =\biggl(\int_{\mathbb{R}^3} f^e_{LR} \frac{1}{q_0}dq\biggl)\biggl(\int_{\mathbb{R}^3} f^e_{LR} dq\int_{\mathbb{R}^3} f^e_{LR} \frac{1}{q^2_0}dq\biggl)^{-\frac{1}{2}}.
\end{align}
Our main result is as follows:
\begin{theorem}\label{main}
Suppose the inflow boundary data $f_{LR}$ is non-negative and belongs to $L^1(\mathbb{R}^3_q)$.
Assume further that $a_l>0$.
Then, we can find $\varepsilon>0$ such that if $w<\varepsilon$, then there exists a unique mild solution $f$ to \eqref{SRBGK} such that
$$
\int_{\mathbb{R}^3} f\frac{1}{q^2_0}dq\geq a_l,\qquad \int_{\mathbb{R}^3} fdq\le a_u, \qquad \frac{K_1}{K_2}(\beta)\le \sqrt{\lambda}.
$$
\end{theorem}
\begin{remark}
(1) $f_{LR}\in L^1(\mathbb{R}^3_q)$ guarantees that $a_l, a_u$ are well-defined, and the condition $a_l>0$ guarantees that the divisor in the ratio $\lambda$ is non-zero.\newline
\noindent (2) Considering the conditions under which the identity holds in the H\"{o}lder inequality, we see that $a_{\ell}>0$ implies $0<\lambda<1$. Therefore, $\sqrt{\lambda}$ is
strictly less than 1 (see the following paragraph below to see why this condition is important).
\end{remark}
To prove our main result, we adapt and make a relativistic extension of the argument in \cite{BY} where one of the authors considered  the stationary problem of the classical ellipsoidal BGK model for classical particles.
The relativistic nature of the equation complicates the problem at virtually every point, and makes the adaptation nontrivial.
One of the key differences arises in  the way the relativistic counterpart of the local temperature $1/\beta$ is defined, which is implicitly defined through a nonlinear functional relation:
\begin{align*}
 \frac{K_1}{K_2}(\beta)=\frac{1}{n}\int_{\mathbb{R}^3}f\frac{1}{q_0}dq.
\end{align*}
This implies that we need to control $(K_1/K_2)(\beta)$  to get a proper bound on $\beta$. 
In view of this, we first  note that we already have some control on it: $0<(K_1/K_2)(\beta)<1$,
which holds trivially by the definitions of $K_1$ and $K_2$.  This trivial bound, however, gives no information on the size of $\beta$ since
$(K_1/K_2)^{-1}(1)=\infty$. And without the information on the size of $\beta$,
we cannot guarantee that our solution space is invariant under our solution map, which is essential to close the
fixed point argument.
Therefore, we need to bound $K_1/K_2$ by a constant that is  strictly less than 1 (see the remark 1.3 (2)). This is accomplished in Lemma \ref{222} using the
following estimate controlling the relativistic Maxwellian by the collision frequency and the boundary data:
 \begin{align*}
&\int_{q_1>0}\frac{w}{|q_1|}\int_0^xe^{-\frac{w}{|q_1|}(x-y)}J_fdy\frac{dq}{q_0}+	\int_{q_1<0}\frac{w}{|q_1|}\int_x^1e^{-\frac{w}{|q_1|}(y-x)}J_fdy\frac{dq}{q_0}\cr
&\hspace{1.4cm}\le \frac{16C_1}{C^2_2}\left(2w\ln\frac{1}{w}+(1+e)w+\frac{\sqrt{2}w^2}{C_2}e^{-\frac{C_2}{\sqrt{2}w}}\right),
\end{align*}
which is established in Lemma \ref{phi2}.\newline


The paper is organized as follows: In Section 2, we define the solution space and present several preliminary technical estimates.
In Section 3, we define the solution operator and  show that it maps the solution space into itself.
In Section 4, we show that the solution operator is a contraction mapping on the solution space under our assumptions.
\newline
%
%
%
%

\section{Estimates in solution space}
We define our solution space $\Omega$ by
$$
\Omega=\left\{f(x,q)\in L^1\left([0,1]\times\mathbb{R}^3\right)~|~ f~ \text{satisfies}~ (\mathcal{A}) \right\}
$$
where the property $(\mathcal{A})$ denotes
$$
\qquad f\ge 0,\qquad a_l\le \int_{\mathbb{R}^3} f\frac{1}{q^2_0}dq,\qquad \int_{\mathbb{R}^3} fdq\le a_u, \qquad \frac{K_1}{K_2}(\beta)\le \sqrt{\lambda}.
$$
We will show that the solution to the boundary value problem (\ref{SRBGK}) is given as a unique fixed point in $\Omega$ of a solution operator, which will be defined later. First, we need to establish several
preliminary estimates.

\begin{lemma}\label{11}
Let $f\in\Omega$, then the macroscopic quantities $n, u, \beta$ constructed from $f$ by the relation (\ref{macro fields}) satisfy
$$
a_l\le n\le a_u,\qquad |u|\le \frac{\sqrt{2}a_u}{a_l},\qquad  \frac{a_l}{a_u}\leq \frac{K_1}{K_2}(\beta)\leq  \sqrt{\lambda}.
$$
\end{lemma}
\begin{proof}First, we see from the definition of $n$ that
$$
n^2=\biggl(\int_{\mathbb{R}^3}f dq\biggl)^2-\sum_{i=1}^3\biggl(\int_{\mathbb{R}^3}f\frac{q_i}{q_0} dq\biggl)^2\le \biggl(\int_{\mathbb{R}^3}f dq\biggl)^2\le a_u^2.
$$
For the lower bound of $n$, we employ the H\"{o}lder inequality as follows:
\begin{align*}
\sum_{i=1}^3\biggl(\int_{\mathbb{R}^3}f\frac{q_i}{q_0} dq\biggl)^2&\leq\sum_{i=1}^3\int_{\mathbb{R}^3}f dq\int_{\mathbb{R}^3}f \frac{q_i^2}{q_0^2}dq\cr
&=\int_{\mathbb{R}^3}f dq\sum_{i=1}^3\int_{\mathbb{R}^3}f \frac{q_i^2}{q_0^2}dq\cr
&=\int_{\mathbb{R}^3}f dq\int_{\mathbb{R}^3}f \frac{|q|^2}{q_0^2}dq.
\end{align*}
to get
\begin{align}\label{holder}
\begin{split}
n^2&=\biggl(\int_{\mathbb{R}^3}f dq\biggl)^2-\sum_{i=1}^3\biggl(\int_{\mathbb{R}^3}f\frac{q_i}{q_0} dq\biggl)^2\cr
&\geq\biggl(\int_{\mathbb{R}^3}f dq\biggl)^2-\int_{\mathbb{R}^3}f dq\int_{\mathbb{R}^3}f \frac{|q|^2}{q_0^2}dq\cr
&=\biggl(\int_{\mathbb{R}^3}f dq\biggl)\biggl(\int_{\mathbb{R}^3}f dq-\int_{\mathbb{R}^3}f \frac{|q|^2}{q_0^2}dq\biggl)\cr
&=\int_{\mathbb{R}^3}f dq\int_{\mathbb{R}^3}f \frac{1}{q^2_0}dq\cr
&\ge \biggl(\int_{\mathbb{R}^3}f \frac{1}{q^2_0}dq\biggl)^2\cr
&= a_l^2.
\end{split}
\end{align}
Using this, we can bound $|u|$ from above:
\[
\displaystyle|u|=\frac{|nu|}{n}\le \frac{1}{a_l}\bigg|\int_{\mathbb{R}^3} f\frac{q}{q_0} dq\bigg|\le \frac{\sqrt{2}}{a_l}\int_{\mathbb{R}^3} f\frac{|q|}{\sqrt{1+|q|^2}} dq\le \frac{\sqrt{2}}{a_l}\int_{\mathbb{R}^3} fdq \le \frac{\sqrt{2}a_u}{a_l},
\]	
and bound $K_1/{K_2}$ from below:
$$
\sqrt{\lambda}\ge\frac{K_1}{K_2}(\beta)=\frac{1}{n}\int_{\mathbb{R}^3}f \frac{1}{q_0}dq\ge \frac{1}{a_u}\int_{\mathbb{R}^3}f \frac{1}{q^2_0}dq\ge \frac{a_l}{a_u}.
$$
\end{proof}	
\begin{lemma}\label{22}
For $f\in \Omega$, there exist constants $C_{1}$ and $C_2$ depending on $a_l,~ a_u$ such that
$$
J_f \le C_{1}e^{-C_{2}\sqrt{1+|q|^2}}.
$$
\end{lemma}
\begin{proof}
It is shown in \cite{BCNS} that $(K_1/K_2)(\beta)$ is strictly increasing, and the range is $[0,1)$ for $\beta\in(0,\infty)$. Therefore, from the result
of Lemma \ref{11}:
$$
\frac{a_l}{a_u}\le\frac{K_1}{K_2}(\beta)\le \sqrt{\lambda},
$$
we can specify the range of $\beta$ as follows:	
\begin{equation}\label{b} \beta_l\equiv\left(\frac{K_1}{K_2}\right)^{-1}\left(\frac{a_l}{a_u}\right)\le\beta
\le\left(\frac{K_1}{K_2}\right)^{-1}\left(\sqrt{\lambda}\right)\equiv\beta_u.
\end{equation}
This, together with  the upper bound of $n$ in Lemma \ref{11} and the fact that $M(\beta)$ is a decreasing function,  gives
\begin{equation}\label{nm}
\frac{n}{M(\beta)}\le \frac{a_u}{M(\beta_u)}.
\end{equation}	
On the other hand, we recall $0\leq |u|\leq \sqrt{2}a_u/a_l$ from Lemma \ref{11}, and use the fact that $h(x)\equiv\sqrt{1+x^2}-x$ is a non-negative decreasing function to
conclude that
\begin{align}\begin{split}\label{uq}
\sqrt{1+|u|^2}\sqrt{1+|q|^2}-u\cdot q&\ge  \sqrt{1+|u|^2}\sqrt{1+|q|^2}-|u||q|\cr
&\ge (\sqrt{1+|u|^2}-|u|)\sqrt{1+|q|^2}\cr
&\ge C_0\sqrt{1+|q|^2},
\end{split}
\end{align}
for
$$
C_0=h\bigg(\frac{\sqrt{2}a_u}{a_l}\bigg)=\sqrt{
	1+\bigg(\frac{\sqrt{2}a_u}{a_l}\bigg)^2}-\frac{\sqrt{2}a_u}{a_l}>0.
$$
Combining \eqref{nm} and \eqref{uq}, we  obtain the desired result:
$$
J_f=\frac{n}{M(\beta)}e^{-\beta\left(\sqrt{1+|u|^2}\sqrt{1+|q|^2}-u\cdot q\right)}\le~ \frac{a_u}{M(\beta_u)}e^{-\beta_lC_0\sqrt{1+|q|^2}}\equiv C_1e^{-C_2\sqrt{1+|q|^2}}.
$$
\end{proof}
\begin{lemma}\label{phi2}
Let $f\in\Omega$. Assume $0<w<1$. Then we have
\begin{align*}
\int_{q_1>0}\frac{w}{|q_1|}\int_0^xe^{-\frac{w}{|q_1|}(x-y)}J_fdydq+	&\int_{q_1<0}\frac{w}{|q_1|}\int_x^1e^{-\frac{w}{|q_1|}(y-x)}J_fdydq\cr
&\le \frac{16C_1}{C^2_2}\left(2w\ln\frac{1}{w}+(1+e)w+\frac{\sqrt{2}w^2}{C_2}e^{-\frac{C_2}{\sqrt{2}w}}\right).
\end{align*}
\end{lemma}
\begin{proof} We only consider $\int_{q_1>0}$ to avoid the repetition. From Lemma \ref{22},
\begin{align}\label{r.h.s}
\begin{split}
\int_{q_1>0}\frac{w}{q_1}\int_0^xe^{-\frac{w}{q_1}(x-y)}J_fdydq&\le C_1\int_{q_1>0}\frac{w}{q_1}\int_0^xe^{-\frac{w}{q_1}(x-y)}e^{-C_2\sqrt{1+|q|^2}}dydq\cr&\le \frac{8C_1}{C^2_2}\int_{q_1>0}\frac{w}{q_1}\int_0^xe^{-\frac{w}{q_1}(x-y)}e^{-\frac{C_2}{\sqrt{2}}|q_1|}dydq_1.
\end{split}
\end{align}
Here we used
\begin{align*}
\int e^{-C_2\sqrt{1+|q|^2}}dq_2dq_3&\le \int e^{-\frac{C_2}{\sqrt{2}}(|q_1|+|q_2|+|q_3|)}dq_2dq_3 \cr
&=e^{-\frac{C_2}{\sqrt{2}}|q_1|}\int e^{-\frac{C_2}{\sqrt{2}}(|q_2|+|q_3|)}dq_2dq_3\cr
&=\frac{8}{C_2^2}e^{-\frac{C_2}{\sqrt{2}}|q_1|}.
\end{align*}
Now we split integral on the r.h.s of (\ref{r.h.s}) into the following two parts:
$$
\int_{q_1>0}=\underbrace{\int_{0<q_1\le\frac{1}{w}}}_{I}+\underbrace{\int_{q_1>\frac{1}{w}}}_{II}.
$$
$\bullet$ (Estimate for $I$): We split $I$ further as
\begin{align*}
I&=\int_{0<q_1\le\frac{1}{w}}e^{-\frac{C_2}{\sqrt{2}}|q_1|}(1-e^{-\frac{w}{q_1}x})dq_1 \cr
&=\biggl\{\underbrace{\int_{0<q_1\le w}}_{I_{1}}+\underbrace{\int_{w<q_1\le \frac{1}{w}}}_{I_{2}}\biggl\}e^{-\frac{C_2}{\sqrt{2}}|q_1|}(1-e^{-\frac{w}{q_1}x})dq_1.
\end{align*}
For $I_{1}$ we have
$$
I_1\le	\int_{0<q_1\le w}1-e^{-\frac{w}{q_1}x}dq_1\le 	\int_{0<q_1\le w}dq_1\le w.
$$
For $I_2$, we use Taylor expansion to estimate
\begin{align*}
I_2&=\int_{w<q_1\le\frac{1}{w}}e^{-\frac{C_2}{\sqrt{2}}|q_1|}\left(\frac{w}{q_1}x-\frac{1}{2!}\left(\frac{w}{q_1}x\right)^2+\frac{1}{3!}\left(\frac{w}{q_1}x\right)^3-\frac{1}{4!}\left(\frac{w}{q_1}x\right)^4+\cdots\right)dq_1\cr
&\le \int_{w<q_1\le\frac{1}{w}}\left(\frac{w}{q_1}+\frac{1}{2!}\left(\frac{w}{q_1}\right)^2+\frac{1}{3!}\left(\frac{w}{q_1}\right)^3+\frac{1}{4!}\left(\frac{w}{q_1}\right)^4+\cdots\right)dq_1\cr
&=2w\ln \frac{1}{w}+\frac{w}{2!}(1-w^2)+\frac{w}{2\cdot3!}(1-w^4)+\frac{w}{3\cdot4!}(1-w^6)+\cdots+\cdots\cr
&\le 2w\ln \frac{1}{w}+w\bigg(1+\frac{1}{2!}+\frac{1}{3!}+\frac{1}{4!}+\cdots\bigg) \cr
&= 2w\ln\frac{1}{w}+we.
\end{align*}
Therefore, we have
\begin{align}\label{I}
I\le 2w\ln\frac{1}{w}+(1+e)w.
\end{align}
$\bullet$ (Estimate for $II$): Since $0\leq x\leq1$, we can bound $II$ as
\begin{align}\label{II}
\begin{split}
II
\le w^2 \int_{q_1>\frac{1}{w}}e^{-\frac{C_2}{\sqrt{2}}|q_1|}dq_1
= \frac{\sqrt{2}w^2}{C_2}e^{-\frac{C_2}{\sqrt{2}w}}.
\end{split}
\end{align}
Combining \eqref{I} and \eqref{II}, we get the desired result:
\begin{equation*}
\int_{q_1>0}\frac{w}{q_1}\int_0^xe^{-\frac{w}{q_1}(x-y)}J_fdydq\le \frac{8C_1}{C^2_2}\left(2w\ln\frac{1}{w}+(1+e)w+\frac{\sqrt{2}w^2}{C_2}e^{-\frac{C_2}{\sqrt{2}w}}\right).
\end{equation*}
\end{proof}
\section{$\Phi$ maps  $\Omega$ into itself}
For $f\in\Omega$, we define our solution operator $\Phi(f)$ as follows:
\begin{align*}
\Phi(f)(x,q)&\equiv\left(e^{-\frac{w}{|q_{1}|}x}f_L+\frac{w}{|q_1|}\int_0^xe^{-\frac{w}{|q_1|}(x-y)}J_fdy\right)1_{q_1>0}\cr
&+\left(e^{-\frac{w}{|q_{1}|}(1-x)}f_R+\frac{w}{|q_1|}\int_x^1e^{-\frac{w}{|q_1|}(y-x)}J_fdy\right)1_{q_1<0}\cr
&\equiv\Phi^+(f)1_{q_1>0}+\Phi^-(f)1_{q_1<0}.
\end{align*}
The main goal of this section is to prove that if $f$ belongs to $\Omega$, $\Phi(f)$ also belongs to $\Omega$:
\begin{proposition}\label{prop1}
The solution operator maps the solution space $\Omega$ into itself.
\end{proposition}
The proof of this proposition is given in the following two lemmas.
\begin{lemma}\label{111} Let $f\in \Omega$. Then we have
$$
\Phi(f)\ge 0.
$$
\end{lemma}
\begin{proof}
From Lemma \ref{11} and \eqref{b}, we see that
$$
\frac{n_f}{M(\beta)}\ge \frac{a_l}{M(\beta_{l})}>0
$$
which gives the positivity of $J_f$. We thus have
\begin{align}\label{phi}
\begin{split}
\Phi(f)&=\left(e^{-\frac{w}{|q_{1}|}x}f_L+\frac{w}{|q_1|}\int_0^xe^{-\frac{w}{|q_1|}(x-y)}J_fdy\right)1_{q_1>0}\cr
&+\left(e^{-\frac{w}{|q_{1}|}(1-x)}f_R+\frac{w}{|q_1|}\int_x^1e^{-\frac{w}{|q_1|}(y-x)}J_fdy\right)1_{q_1<0}\cr
&\ge e^{-\frac{w}{|q_{1}|}x}f_L1_{q_1>0}+e^{-\frac{w}{|q_{1}|}(1-x)}f_R1_{q_1<0}\cr
&\ge 0.
\end{split}
\end{align}
\end{proof}

\begin{lemma}\label{222} Let $f\in\Omega$. Then, for sufficiently small $w$, $\Phi(f)$ satisfies
$$
\int_{\mathbb{R}^3} \Phi(f)\frac{1}{q^2_0}dq\ge a _l,\qquad \int_{\mathbb{R}^3}\Phi(f)dq\le a_u,\qquad 	\frac{1}{n_{\Phi}}\int_{\mathbb{R}^3} \Phi(f)\frac{1}{q_0}dq\le \sqrt{\lambda}
$$
where $n_\Phi$ denotes the proper particle density with respect to $\Phi$:
$$
n_\Phi^2=\biggl(\int_{\mathbb{R}^3}\Phi(f) dq\biggl)^2-\sum_{i=1}^3\biggl(\int_{\mathbb{R}^3}\Phi(f)\frac{q_i}{q_0} dq\biggl)^2.
$$
\end{lemma}
\begin{proof}
(1) We have from \eqref{phi}
\begin{align*}
\Phi(f)\frac{1}{q^2_0}&\ge e^{-\frac{w}{|q_{1}|}x}f_L\frac{1}{q^2_0}~1_{q_1>0}+e^{-\frac{w}{|q_{1}|}(1-x)}f_R\frac{1}{q^2_0}~1_{q_1<0}\cr
&\ge e^{-\frac{w}{|q_{1}|}}f_L\frac{1}{q^2_0}~1_{q_1>0}+e^{-\frac{w}{|q_{1}|}}f_R\frac{1}{q^2_0}~1_{q_1<0}\cr
&= e^{-\frac{w}{|q_{1}|}}f_{LR}\frac{1}{q^2_0},
\end{align*}
yielding
$$
\int_{\mathbb{R}^3} \Phi(f)\frac{1}{q^2_0}dq\ge \int_{\mathbb{R}^3} e^{-\frac{w}{|q_{1}|}}f_{LR}\frac{1}{q^2_0}dq =a_l.
$$	
(2) We recall Lemma \ref{phi2} to see
\begin{align*}
\int_{\mathbb{R}^3}\Phi(f)dq&=\int_{\mathbb{R}^3} e^{-\frac{w}{|q_{1}|}x}f_L1_{q_1>0}+e^{-\frac{w}{|q_{1}|}(1-x)}f_R1_{q_1<0}dq\cr
&+\int _{\mathbb{R}^3}\frac{w}{|q_1|}\int_0^xe^{-\frac{w}{|q_1|}(x-y)}J_fdy1_{q_1>0}+\frac{w}{|q_1|}\int_x^1e^{-\frac{w}{|q_1|}(y-x)}J_fdy1_{q_1<0}dq\cr
&\le\int_{\mathbb{R}^3} f_{LR}dq+\frac{16C_1}{C^2_2}\left(2w\ln\frac{1}{w}+(1+e)w+\frac{\sqrt{2}w^2}{C_2}e^{-\frac{C_2}{\sqrt{2}w}}\right).
\end{align*}
We then choose a sufficiently small $w$  so that
$$
\frac{16C_1}{C^2_2}\left(2w\ln\frac{1}{w}+(1+e)w+\frac{\sqrt{2}w^2}{C_2}e^{-\frac{C_2}{\sqrt{2}w}}\right)\le \int_{\mathbb{R}^3} f_{LR} dq
$$
to get
$$
\int_{\mathbb{R}^3} \Phi(f)dq\le 2\int_{\mathbb{R}^3} f_{LR} dq=a_u.
$$
(3) Estimating similarly as in \eqref{holder}, we get
\begin{align*}
n_{\Phi}&\geq\biggl(\int_{\mathbb{R}^3}\Phi(f) dq\int_{\mathbb{R}^3}\Phi(f) \frac{1}{q^2_0}dq\biggl)^{\frac{1}{2}}.
\end{align*}
But we have from the positivity of $J_f$
\begin{align*}
f^e_{LR}(x,q)&=e^{-\frac{w}{|q_{1}|}x}f_L1_{q_1>0}+e^{-\frac{w}{|q_{1}|}(1-x)}f_R1_{q_1<0}\cr
&\le \left(e^{-\frac{w}{|q_{1}|}x}f_L+\frac{w}{|q_1|}\int_0^xe^{-\frac{w}{|q_1|}(x-y)}J_fdy\right)1_{q_1>0}\cr
&+\left(e^{-\frac{w}{|q_{1}|}(1-x)}f_R+\frac{w}{|q_1|}\int_x^1e^{-\frac{w}{|q_1|}(y-x)}J_fdy\right)1_{q_1<0}\cr
&=\Phi(f)(x,q),
\end{align*}
so that we can bound $n_{\Phi}$ further from below as
\begin{align*}
n_{\Phi}\geq\biggl(\int_{\mathbb{R}^3} f^e_{LR}dq\int_{\mathbb{R}^3} f^e_{LR}\frac{1}{q^2_0}dq\biggl)^{\frac{1}{2}}.
\end{align*}
Therefore, we get
\begin{align*}
\begin{split}
\frac{1}{n_{\Phi}}\int_{\mathbb{R}^3} \Phi(f)\frac{1}{q_0}dq&\leq\biggl(\int_{\mathbb{R}^3}\Phi(f) dq\int_{\mathbb{R}^3}\Phi(f) \frac{1}{q^2_0}dq\biggl)^{-\frac{1}{2}}\int_{\mathbb{R}^3} \Phi(f)\frac{1}{q_0}dq\cr
&\leq\biggl(\int_{\mathbb{R}^3} f^e_{LR}dq\int_{\mathbb{R}^3} f^e_{LR}\frac{1}{q^2_0}dq\biggl)^{-\frac{1}{2}}\int_{\mathbb{R}^3} \Phi(f)\frac{1}{q_0}dq,
\end{split}
\end{align*}
which, in view of the definition of $\lambda$ in (\ref{lambda}), leads to
\begin{align}\label{phi6}
\begin{split}
\frac{1}{n_{\Phi}}\int_{\mathbb{R}^3} \Phi(f)\frac{1}{q_0}dq
&\leq \lambda\biggl(\int_{\mathbb{R}^3} f^e_{LR}\frac{1}{q_0}dq\biggl)^{-1}\int_{\mathbb{R}^3} \Phi(f)\frac{1}{q_0}dq.
\end{split}
\end{align}
Now, since
\begin{align*}
\int_{\mathbb{R}^3}\Phi(f)\frac{1}{q_0}dq&=\int_{\mathbb{R}^3} f^e_{LR} \frac{1}{q_0}dq+\int_{q_1>0}\frac{w}{|q_1|}\int_0^xe^{-\frac{w}{|q_1|}(x-y)}J_fdy\frac{1}{q_0}dq\cr
&+\int_{q_1<0}\frac{w}{|q_1|}\int_x^1e^{-\frac{w}{|q_1|}(y-x)}J_fdy\frac{1}{q_0}dq,
\end{align*}
Lemma \ref{phi2} implies
\begin{align*}
\int_{\mathbb{R}^3}\Phi(f)\frac{1}{q_0}dq\leq\int_{\mathbb{R}^3} f^e_{LR} \frac{1}{q_0}dq+\frac{16C_1}{C^2_2}\left(2w\ln\frac{1}{w}+(1+e)w+\frac{\sqrt{2}w^2}{C_2}e^{-\frac{C_2}{\sqrt{2}w}}\right).
\end{align*}
We then note that, as $w$ decreases,
$$
f^e_{LR}=e^{-\frac{w}{|q_{1}|}x}f_L1_{q_1>0}+e^{-\frac{w}{|q_{1}|}(1-x)}f_R1_{q_1<0}
$$
increases, which enables one to find $w$ sufficiently small such that
$$
\frac{16C_1}{C^2_2}\left(2w\ln\frac{1}{w}+(1+e)w+\frac{\sqrt{2}w^2}{C_2}e^{-\frac{C_2}{\sqrt{2}w}}\right)\le \left(\frac{1}{\sqrt{\lambda}}-1 \right)\int_{\mathbb{R}^3} f^e_{LR} \frac{1}{q_0}dq,
$$
yielding
\begin{align*}
\begin{split}
\int_{\mathbb{R}^3}\Phi(f)\frac{1}{q_0}dq&\le\int_{\mathbb{R}^3} f^e_{LR} \frac{1}{q_0}dq+\frac{16C_1}{C^2_2}\left(2w\ln\frac{1}{w}+(1+e)w+\frac{\sqrt{2}w^2}{C_2}e^{-\frac{C_2}{\sqrt{2}w}}\right)\cr
&\le \frac{1}{\sqrt{\lambda}}\int_{\mathbb{R}^3} f^e_{LR} \frac{1}{q_0}dq.
\end{split}
\end{align*}
Inserting this into \eqref{phi6} gives the desired result.
\end{proof}

\section{Contraction mapping}
In this section, we establish the Lipschitz continuity of our solution operator. We first need to set up preliminary
computations. The following lemma can be found  in \cite{BCNS}, but we provide a detailed proof for reader's convenience.
\begin{lemma}\label{bessel} For the modified Bessel function of the second kind $K_i(\beta)$, the following holds
$$
\biggl(\frac{K_{1}}{K_{2}}\biggl)^{\prime}(\beta) =\frac{3}{\beta}\frac{K_{1}}{K_{2}}(\beta)+\biggl(\frac{K_{1}}{K_{2}} \biggl)^{2}(\beta)-1.
$$
	
\end{lemma}
\begin{proof}
We recall that
$$
K_i(\beta)=\int_0^\infty \cosh(ir)e^{-\beta\cosh(r)}dr,
$$
 and use change of variable $x=\sinh r$ to get
\begin{align*}
K_{0}(\beta)&=\int_{0}^{\infty}\exp\biggl\{-\beta
\cosh(r)\biggl\}dr=\int_{0}^{\infty}\frac{1}{\sqrt{1+x^{2}}}\exp\biggl\{-\beta
\sqrt{1+x^{2}}\biggl\}dx,\cr
K_{1}(\beta)&=\int_{0}^{\infty}\cosh(r)\exp\biggl\{-\beta \cosh(r)\biggl\}dr=\int_{0}^{\infty}\exp\biggl\{-\beta \sqrt{1+x^{2}}\biggl\}dx, \cr
K_{2}(\beta)&=\int_{0}^{\infty}\cosh(2r)\exp\biggl\{-\beta
\cosh(r)\biggl\}dr=\int_{0}^{\infty}\frac{2x^{2}+1}{\sqrt{1+x^{2}}}\exp\biggl\{-\beta
\sqrt{1+x^{2}}\biggl\}dx.
\end{align*}
We then observe that
\begin{align*}
(K_{1})^{\prime}(\beta)&=-\int_{0}^{\infty}\sqrt{1+x^{2}}\exp\biggl\{-\beta \sqrt{1+x^{2}}\biggl\}dx\cr
&=-\frac{1}{2}\biggl(K_{2}(\beta)+K_{0}(\beta)\biggl)\cr
&=-\biggl(\frac{1}{\beta}K_{1}(\beta)+K_{0}(\beta)\biggl)
\end{align*}
and
\begin{align*}
(K_{2})^{\prime}(\beta)&=-\int_{0}^{\infty}(2x^{2}+1)\exp\biggl\{-\beta \sqrt{1+x^{2}}\biggl\}dx\cr
&=\frac{2}{\beta}\int_{0}^{\infty}x\sqrt{1+x^{2}}\frac{d}{dx}\biggl(\exp\biggl\{-\beta \sqrt{1+x^{2}}\biggl\}\biggl)dx-K_{1}(\beta)\cr
&=-\frac{2}{\beta}\int_{0}^{\infty}\frac{2x^{2}+1}{\sqrt{1+x^{2}}}\exp\biggl\{-\beta \sqrt{1+x^{2}}\biggl\}dx-K_{1}(\beta)\cr
&=-\frac{2}{\beta}K_{2}(\beta)-K_{1}(\beta)\cr
&=\biggl(-\frac{4}{\beta^{2}}-1\biggl)K_{1}(\beta)-\frac{2}{\beta}K_{0}(\beta).
\end{align*}
Here we used
\begin{align*}
K_{2}(\beta)&=\int_{0}^{\infty}\frac{2x^{2}+1}{\sqrt{1+x^{2}}}\exp\biggl\{-\beta \sqrt{1+x^{2}}\biggl\}dx\cr
&=\int_{0}^{\infty}\frac{2x^{2}}{\sqrt{1+x^{2}}}\exp\biggl\{-\beta \sqrt{1+x^{2}}\biggl\}dx+K_{0}(\beta)\cr
&=\int_{0}^{\infty}-\frac{2x}{\beta}\frac{d}{dx}\biggl(\exp\biggl\{-\beta \sqrt{1+x^{2}}\biggl\}\biggl)dx+K_{0}(\beta)\cr
&=\frac{2}{\beta}\int_{0}^{\infty}\exp\biggl\{-\beta \sqrt{1+x^{2}}\biggl\}dx+K_{0}(\beta)\cr
&=\frac{2}{\beta}K_{1}(\beta)+K_{0}(\beta).
\end{align*}
These identities then give
\begin{equation*}
(K_{1})^{\prime}(\beta)K_{2}(\beta)-K_{1}(\beta)(K_{2})^{\prime}(\beta)=\frac{3}{\beta}K_{1}(\beta)K_{2}(\beta)+(K_{1})^{2}(\beta)-(K_{2})^{2}(\beta),
\end{equation*}	
which, upon dividing both sides by  $(K_{2})^2(\beta)$, gives the desired result.
\end{proof}
The following lemma shows that the r.h.s of the identity in Lemma \ref{bessel} is strictly positive. The proof can be
found in \cite{BCNS}.
\begin{lemma}\cite{BCNS}\label{bessel2} For the modified Bessel function of the second kind, the following inequality holds
\begin{align*}
		\frac{3}{\beta}\frac{K_{1}}{K_{2}}(\beta)+\biggl(\frac{K_{1}}{K_{2}}\biggl)^{2}(\beta)-1
		\geq\ell(\beta),
	\end{align*}
	where $\ell(\beta)$ is defined by
	\begin{align*}
		\ell(\beta)\equiv\begin{cases}\frac{2-\beta}{(\beta+2)^{2}} & 0<\beta<2\cr \frac{3(6656\beta^{4}+2419\beta^{3}+726)}{(128\beta^{3}+240\beta^{2}+105\beta-66)^{2}}&\beta\geq2.\end{cases}
	\end{align*}
Note that $\ell(\beta)$ is strictly positive.
\end{lemma}
\begin{lemma}\label{JJ} Let $f,g\in\Omega$, then, for sufficiently small $w$, there exist positive constants $C_8$ and $C_9$ such that
$$
|J_f-J_g|\le C_9e^{-C_8\sqrt{1+|q|^2}}\|f-g\|_{L^1_q},
$$
where $\|\cdot\|_{L^1_q}$ denotes the usual $L^1$ norm:
\[
\|f\|_{L^1_q}=\int_{\mathbb{R}^{3}}|f(q)|dq.
\]
\end{lemma}
\begin{proof}
For the convenience of computation, we introduce a new variable $\alpha$ (see \cite{BCNS}) defined by
$$
\alpha =\frac{1}{n}\int_{\mathbb{R}^3}f\frac{1}{q_0}dq.
$$
Due to the monotonicity of $K_1/K_2$ established in Lemma \ref{bessel} and Lemma \ref{bessel2}, there is a one-to-one correspondence between $\alpha$ and $\beta$:
\[
\beta=\mathcal{X}(\alpha)=\biggl( \frac{K_1}{K_2}\biggl)^{-1}(\alpha).
\]
In view of this, we consider  $J(n,u,\beta)$ as a functional of $(n,u,\alpha)$, and apply the mean value theorem to get
\begin{align}\label{J}
J(n_f,u_f,\alpha_f)-J(n_g,u_g,\alpha_g)=\nabla_{n,u,\alpha} J(\theta)\cdot\big(n_f-n_g,u_f-u_g,\alpha_f-\alpha_g\big),
\end{align}	
for some $0\leq \theta\leq 1$, where the abbreviate notation $J(\theta)$ denotes
\[
J(\theta)=J\big((1-\theta)n_f+\theta n_g,(1-\theta)u_{f}+\theta u_g,(1-\theta)\alpha_{f}+\theta \alpha_g\big).
\]
We need to estimate $\nabla_{n,u,\alpha} J$ and $(n_f-n_g,u_f-u_g,\alpha_f-\alpha_g\big)$. \newline

\noindent(1) Estimates for $\nabla_{n,u,\alpha} J$: A direct computation gives
\begin{align}\label{return}
\begin{split}
&\frac{\partial J}{\partial n}=\frac{1}{n}J,\qquad  		 \nabla_u J=\beta\biggl(q-\frac{\sqrt{1+|q|^2}}{\sqrt{1+|u|^2}}u\biggl)J,\cr
&\frac{\partial J}{\partial \alpha}=-\frac{\partial\beta}{\partial\alpha}\biggl(\frac{M^{\prime}(\beta)}{M(\beta)}+\sqrt{1+|q|^2}\sqrt{1+|u|^2}-u\cdot q\biggl)J.
\end{split}
\end{align}
Using Lemma \ref{11}, Lemma \ref{22} and \eqref{b}, we can show that $\frac{\partial J}{\partial n}$ and $\nabla_u J$ are bounded as
\begin{align*}
\left|\frac{\partial J}{\partial n}\right|&\le \frac{C_{1}}{a_l}e^{-C_{2}\sqrt{1+|q|^2}},\cr
\left|\nabla_u J\right|&\le\beta_u\biggl(|q|+\sqrt{1+|q|^2}\frac{|u|}{\sqrt{1+|u|^2}}\biggl)|J| \cr
&\le 2C_{1}\beta_u\sqrt{1+|q|^2}e^{-C_{2}\sqrt{1+|q|^2}}\cr
&\le 2C_{1}\beta_uC_3e^{-C_4\sqrt{1+|q|^2}}.
\end{align*}
The estimate for $	\frac{\partial }{\partial \alpha}J$ is more involved. First, we use differentiation rule for inverse functions and Lemma \ref{bessel} to get
\begin{align*}
\frac{\partial \beta}{\partial \alpha}&=\frac{\partial}{\partial\alpha}	\biggl(\frac{K_{1}}{K_{2}}\biggl)^{-1}(\alpha)
=\frac{1}{\big(\frac{K_{1}}{K_{2}}\big)^{\prime}(\beta)}
=\frac{1}{\frac{3}{\beta}\frac{K_{1}}{K_{2}}(\beta)+\big(\frac{K_{1}}{K_{2}}\big)^{2}(\beta)-1}.
\end{align*}
We then recall Lemma \ref{bessel2} that
\begin{align*}
\frac{3}{\beta}\frac{K_{1}}{K_{2}}(\beta)+\biggl(\frac{K_{1}}{K_{2}}\biggl)^{2}(\beta)-1
\geq\ell(\beta).
\end{align*}
Note that $\ell(\beta)$ is a strictly positive function. Therefore, we can conclude that the continuous function $\big(\frac{K_{1}}{K_{2}}\big)^{\prime}(\beta)$
(since the modified Bessel function of the second kind is continuous)  is strictly positive
on a closed and bounded interval $[\beta_l,\beta_u]$. This implies that $\big(\frac{K_{1}}{K_{2}}\big)^{\prime}(\beta)$ posses a strictly positive minimum on $[\beta_l,\beta_u]$.
If we denote it by $1/C_5>0$, we have
\begin{align*}
\frac{3}{\beta}\frac{K_{1}}{K_{2}}(\beta)+\biggl(\frac{K_{1}}{K_{2}}\biggl)^{2}(\beta)-1\geq \frac{1}{C_5}.
\end{align*}
In conclusion, we obtain
\begin{align*}
\Big|\frac{\partial\beta}{\partial\alpha}\Big|&=\Big|\frac{3}{\beta}\frac{K_{1}}{K_{2}}(\beta)+\biggl(\frac{K_{1}}{K_{2}}\biggl)^{2}(\beta)-1\Big|^{-1}\cr
&=\Big(\frac{3}{\beta}\frac{K_{1}}{K_{2}}(\beta)+\biggl(\frac{K_{1}}{K_{2}}\biggl)^{2}(\beta)-1\Big)^{-1}\cr
&\leq C_5.
\end{align*}
On the other hand, it is clear that there exists a constant $C_6>0$ such that
$$
\frac{M^{\prime}(\beta)}{M(\beta)}=-\frac{\int_{\mathbb{R}^3} \sqrt{1+|q|^2}\exp\{-\beta\sqrt{1+|q|^2}\}dq}{\int_{\mathbb{R}^3} \exp\{-\beta\sqrt{1+|q|^2}\}dq}<C_6,
$$
when $\beta$ lies in a closed and bounded range: $\beta\in[\beta_l,\beta_u]$. We return back to (\ref{return}) with these estimates to get
\begin{align*}
\left| \frac{\partial J}{\partial \alpha}\right|&\le C_5\left(C_6+2\sqrt{1+|q|^2}\sqrt{1+4a^2_u/a^2_l}\right)C_1e^{-C_2\sqrt{1+|q|^2}}\cr
&\le C_5C_6C_1e^{-C_2\sqrt{1+|q|^2}}+2C_5C_3\sqrt{1+4a^2_u/a^2_l}C_1e^{-C_4\sqrt{1+|q|^2}}\cr
&\le C_7e^{-C_8\sqrt{1+|q|^2}}
\end{align*}
(2) Estimates on $(n_f-n_g ,u_f-u_g ,\alpha_f-\alpha_g )$ \newline
$\bullet~n_f-n_g$: From Lemma \ref{11}, we get
\begin{align*}
|n_{f}-n_{g}|&=\left|\frac{n^2_f-n^2_g}{n_{f}+n_{g}}\right|\cr
&\le\frac{1}{n_{f}+n_{g}}\biggl\{\left(\int_{\mathbb{R}^3}|f-g| dq\right)\left(\int_{\mathbb{R}^3}|f+g| dq\right)\cr
&~-\sum_{i=1}^3\left(\int_{\mathbb{R}^3}|f-g|\frac{q_i}{q_0} dq\right)\left(\int_{\mathbb{R}^3}|f+g|\frac{q_i}{q_0} dq\right)\biggl\}\cr
&\le \frac{1}{2a_l}\biggl(2a_u\int_{\mathbb{R}^3}|f-g| dq+\sum_{i=1}^32a_u\int_{\mathbb{R}^3}|f-g|dq\biggl)\cr
&\le \frac{4a_u}{a_l}\|f-g\|_{L^1_q}.
\end{align*}
$\bullet~u_f-u_g$: Using the above estimate, Lemma \ref{11} and
\begin{align*}
|n_fu_f-n_gu_g|= \left|\int_{\mathbb{R}^3}(f-g)\frac{q}{q_0}dq\right|\le\sqrt{2} \int_{\mathbb{R}^3}|f-g|dq= \sqrt{2}\|f-g\|_{L^1_q},
\end{align*}
we compute
\begin{align*}
|u_{f}-u_{g}|&=\left|\frac{n_g(n_fu_f)-n_f(n_gu_g)}{n_{f}n_{g}}\right|\cr
&=\left|\frac{n_g(n_fu_f-n_gu_g)-n_gu_g(n_f-n_g)}{n_{f}n_{g}}\right|\cr
&\le\frac{\sqrt{2}}{a^2_l}\big(a_u \|f-g\|_{L^1_q}+\sqrt{2}a_u\frac{4a_u}{a_l}\|f-g\|_{L^1_q}\big) \cr
&\le\left( \frac{\sqrt{2}a_u}{a_l^2}+\frac{8a^2_u}{a^3_l}\right)\|f-g\|_{L^1_q}.
\end{align*}
$\bullet~\alpha_f-\alpha_g$:  We can estimate similarly as in the previous cases:
\begin{align*}		
|\alpha_{f}-\alpha_{g}|&=\frac{1}{n_fn_g}\biggl|n_g\int_{\mathbb{R}^3}f\frac{1}{q_0}dq-n_f\int_{\mathbb{R}^3}g\frac{1}{q_0}dq\biggl|\cr
&=\frac{1}{n_fn_g}\biggl|n_g\int_{\mathbb{R}^3}(f-g)\frac{1}{q_0}dq-(n_f-n_g)\int_{\mathbb{R}^3}g\frac{1}{q_0}dq\biggl|\cr
&\le \frac{1}{a^2_l}\big(a_u\|f-g\|_{L^1_q}+a_u\frac{4a_u}{a_l}\|f-g\|_{L^1_q}\big)\cr
&\le \left(\frac{a_u}{a_l^2}+\frac{4a^2_u}{a^3_l}\right)\|f-g\|_{L^1_q}.
\end{align*}
Combining all the estimates we obtained so far, we get
\begin{align*}
\left|\frac{\partial J}{\partial n}(n_f-n_g)\right|&\le \frac{4a_uC_{1}}{a^2_l}e^{-C_{8}\sqrt{1+|q|^2}}\|f-g\|_{L^1_q},\cr
\left|\nabla_u J\cdot (u_{f}-u_{g}) \right|&\le 2C_{1}\beta_u C_3\left( \frac{\sqrt{2}a_u}{a_l^2}+\frac{8a^2_u}{a^3_l}\right)e^{-C_{8}\sqrt{1+|q|^2}}\|f-g\|_{L^1_q},\cr
\left|\frac{\partial J}{\partial \alpha} (\alpha_{f}-\alpha_{g})\right|&\le C_7\left(\frac{a_u}{a_l^2}+\frac{4a^2_u}{a^3_l}\right)e^{-C_8\sqrt{1+|q|^2}}\|f-g\|_{L^1_q}.
\end{align*}
Thus \eqref{J} can be estimates as
\begin{align*}
|J(n_f,u_f,\alpha_f)-J(n_g,u_g,\alpha_g)|&\le  C_9e^{-C_8\sqrt{1+|q|^2}}\|f-g\|_{L^1_q},
\end{align*}
where the constant $C_9$ is given by
\begin{align*}
	C_9&= \frac{a_u}{a^2_l}\left(4C_1+  2C_{1}\beta_u C_3\left( \sqrt{2}+\frac{8a_u}{a_l}\right)+C_7\left(1+\frac{4a_u}{a_l}\right)\right)
\end{align*}
This gives the desired result.
\end{proof}
The following proposition, together with Proposition \ref{prop1} completes the proof of Theorem \ref{main}.
\begin{proposition} $\Phi(f)$ is a contraction mapping on $\Omega$ for sufficiently small $w$. That is, we can take
$w$ sufficiently small such that there exists $0<\alpha<1$ satisfying
\[
\sup_x\|\Phi(f)-\Phi(g)\|_{L^1_q}\leq\alpha \sup_x\|f-g\|_{L^1_q}.
\]
\end{proposition}
\begin{proof}
We have for $f,g\in \Omega$
\begin{align*}
&\int_{\mathbb{R}^3} |\Phi(f)-\Phi(g)|dq\cr
&\qquad=\int_{q_1>0}|\Phi(f)-\Phi(g)|dq+\int_{q_1<0}|\Phi(f)-\Phi(g)|dq \cr
&\qquad\leq\int_{q_1>0}\frac{w}{q_1}\int_0^xe^{-\frac{w}{q_1}(x-y)}|J_f-J_g|dydq+\int_{q_1<0} \frac{w}{|q_1|}\int_x^1e^{-\frac{w}{|q_1|}(y-x)}|J_f-J_g|dydq.
\end{align*}
Then, using Lemma \ref{phi2} and Lemma \ref{JJ}, we can control the last term as follows:
\begin{align*}
& \int_{q_1>0}\frac{w}{q_1}\int_0^xe^{-\frac{w}{q_1}(x-y)}|J_f-J_g|dydq+\int_{q_1<0} \frac{w}{|q_1|}\int_x^1e^{-\frac{w}{|q_1|}(y-x)}|J_f-J_g|dydq\cr
&\hspace{1.2cm}\le C_9\biggl\{\int_{q_1>0}\frac{w}{q_1}e^{-C_8\sqrt{1+|q|^2}}\int_0^xe^{-\frac{w}{q_1}(x-y)}dydq\cr
&\hspace{1.2cm}+\int_{q_1<0} \frac{w}{|q_1|}e^{-C_8\sqrt{1+|q|^2}}\int_x^1e^{-\frac{w}{|q_1|}(y-x)}dydq\biggl\}\|f-g\|_{L^1_q}\cr
&\hspace{1.2cm}\le \frac{16C_9}{C^2_8}\left(2w\ln\frac{1}{w}+(1+e)w+\frac{\sqrt{2}w^2}{C_8}e^{-\frac{C_8}{\sqrt{2}w}}\right)\|f-g\|_{L^1_q}.
\end{align*}
For sufficiently small $w$, this gives the desired results.
\end{proof}
\noindent{\bf Acknowledgement}
This research was supported by Basic Science Research Program through the National Research Foundation of Korea(NRF) funded by the Ministry of Education(NRF-2016R1D1A1B03935955)

\end{document}